Check that T = R^3 to R^3, defined by


T(x1,x2,X3)= (x1+X3, x2+2x3, x1-x2-x3) is a linear operator. Also find the kernel

1
Expert's answer
2021-12-06T16:24:33-0500

properties of linear operator:

"T(x+y)=T(x)+T(y)"

"T(cx)=cT(x)"


we have:

"T((x_1,x_2,x_3)+(y_1,y_2,y_3))="

"=(x_1+x_3+y_1+y_3, x_2+2x_3+y_2+2y_3, x_1-x_2-x_3+y_1-y_2-y_3)"

"T(x_1,x_2,x_3)+T(y_1,y_2,y_3)="

"=(x_1+x_3, x_2+2x_3, x_1-x_2-x_3) +(y_1+y_3, y_2+2y_3, y_1-y_2-y_3)="

"=(x_1+x_3+y_1+y_3, x_2+2x_3+y_2+2y_3, x_1-x_2-x_3+y_1-y_2-y_3)"

"T((x_1,x_2,x_3)+(y_1,y_2,y_3))=T(x_1,x_2,x_3)+T(y_1,y_2,y_3)"


"T(cx_1,cx_2,cx_3)=(c(x_1+x_3), c(x_2+2x_3),c( x_1-x_2-x_3))"

"cT(x_1,x_2,x_3)=c(x_1+x_3, x_2+2x_3, x_1-x_2-x_3)="

"=(c(x_1+x_3), c(x_2+2x_3),c( x_1-x_2-x_3))"

"T(cx_1,cx_2,cx_3)=cT(x_1,x_2,x_3)"


so, "T(x_1,x_2,x_3)" is linear operator


for kernel:

"T(x_1,x_2,x_3)=(x_1+x_3, x_2+2x_3, x_1-x_2-x_3) =0"


"x_1+x_3=0"

"x_2+2x_3=0"

"x_1-x_2-x_3=0"

"x_1=-x_3,x_2=-2x_3"


kernel:

"(x_1,x_2,x_3)=(-x_3,-2x_3,x_3)"